ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

8. Analysis

1991 MS Classification: 30 Näätänen, Marjatta, University of Helsinki, Finland Arithmetic Fuchsian groups of signature (0; a, b, c, oo) Using an equation for the corresponding Teichmuller space, arithmetic Fuchsian groups of signature (0; a, b, c, oo) are classified (up conjugacy in PGL(2, R)) in terms of the signature and the describing triple. Joint work with Toshihiro Nakanishi and Gerhard Rosenberger.